OpenStudy (anonymous):

can you solve the square root of 6x+1=x-1

OpenStudy (anonymous):

square both sides. square root of 6x+1^2=(x-1)^2 6x+1=(x-1)^2 6x+1=x2-2x+1 (using foil) x^2-8x=0 x(x-8) x=8